I am not sure if I am doing this problem correctly could I please have some help. The additive inverse of a number divided by twelve is the same as one less than three times its reciprocal. Find the number. I have x+2x=3/10 x=1/10
Thank you.

Answer by scott8148(6628)

"The additive inverse of a number divided by twelve" __ -x/12

"is the same as" __ =

"one less than three times its reciprocal" __ (3/x)-1

-x/12=(3/x)-1 __ multiplying by 12x __ -x^2=36-12x __ adding x^2 __ 0=x^2-12x+36

factoring __ 0=(x-6)(x-6)